# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v2

@@ Abstract
Computing the polar decomposition and the related matrix sign function has been a well-studied problem in numerical analysis for decades. Recently, it has emerged as an important subroutine within the Muon algorithm for training deep neural networks. However, the requirements of this application differ sharply from classical settings: deep learning demands GPU-friendly algorithms that prioritize high throughput over high precision. We introduce Polar Express, a new method for computing the polar decomposition. Like Newton–Schulz and other classical polynomial methods, our approach uses only matrix-matrix multiplications, making it very efficient on GPUs. Inspired by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the update rule at each iteration by solving a minimax optimization problem. We prove that this strategy minimizes error in a worst-case sense, allowing Polar Express to converge as rapidly as possible both in the early iterations and asymptotically. We also address finite-precision issues, making it practical to use in bfloat16. When integrated into the Muon training framework, our method leads to consistent improvements in validation loss when training a GPT-2 model on one billion tokens from the FineWeb dataset, outperforming recent alternatives across a range of learning rates.
@@ 1   Introduction
X = X / (X.norm(dim=(-2, -1), keepdim=True) * 1.01)
@@
Figure 1: Training a GPT-2-Large model (774M params) on 1 billion tokens from the FineWeb dataset [2]. The label muon-<name> refers to implementing Muon using <name> to compute the polar factor. Left: final validation loss across learning rates. Right: validation loss across epochs using the best learning rate. The best learning rate ($lr$) and final validation loss for each method was adamw $(lr=0.0001)$: $4.172$, muon-You $(lr=0.02)$: $3.400$, muon-Jordan $(lr=0.02)$: $3.398$ and muon-PolarExp $(lr=0.02)$: $3.340$.
@@ 1.1   The Muon Method
The Muon optimizer has recently gained popularity for training large language models, often outperforming state-of-the-art adaptive gradient methods like Adam and AdamW [26, 31]. Muon has been used to set records for the NanoGPT speedrun [22] and to expand the Pareto frontier of performance versus training FLOPs for large language models [30, 42].
@@
Whereas standard stochastic gradient descent (SGD) with momentum updates the weight matrix by taking a step in the direction $-\bm{M}_{t}$, the Muon method steps in the direction $-\operatorname*{polar}(\bm{M}_{t})$, where $\operatorname*{polar}(\bm{M})$ denotes the closest semi-orthogonal matrix to $\bm{M}$ [19, Chapter 8]. Concretely, if $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ is the singular value decomposition (SVD) of $\bm{M}$, then
@@
The matrix $\operatorname*{polar}(\bm{M})$ can be seen as a generalization of the matrix sign function to rectangular matrices [3]. Indeed, when $\bm{M}$ is square symmetric with eigendecomposition $\bm{M}=\bm{V}\bm{\Lambda}\bm{V}^{\mathsf{T}}$, $\operatorname*{polar}(\bm{M})$ exactly coincides with the matrix sign function $\operatorname*{sign}(\bm{M})=\bm{V}\operatorname*{sign}(\bm{\Lambda})\bm{V}^{\mathsf{T}}$ [19, Chapter 5]. Equivalently, $\operatorname*{polar}(\bm{M})$ is the left orthogonal factor of the polar decomposition of $\bm{M}$ [19, Chapter 8]. The motivation for Muon is that $-\operatorname*{polar}(\bm{M})$ gives the steepest-descent direction with respect to the spectral norm (instead of the Frobenius norm, as in standard SGD).
Recent work [40] shows that Muon can be viewed as a conditional gradient (Frank-Wolfe) method with a trust region defined by the spectral norm. In the same work, the authors also provide a convergence theory for the smooth and non-convex setting, as well as for the stochastic non-convex case. The analysis of Muon was further refined in [41], which proves convergence under a layerwise $(L_{0},L_{1})$-smoothness assumption, in both the stochastic non-convex and stochastic Polyak–Łojasiewicz settings. We refer the reader to [22] and [5] for further background. In this paper, we take the Muon update rule as given and focus on the problem of efficiently computing the polar decomposition $\operatorname*{polar}(\bm{M})$.
@@ 1.2   Computing the Polar Factor
Although $\operatorname*{polar}(\bm{M})$ can be computed directly via an SVD in $O(\min(mn^{2},nm^{2}))$ time, doing so is prohibitively expensive in deep learning applications, especially as standard SVD algorithms fail to take full advantage of the parallelism available on GPUs. There has been significant work on highly-parallel methods for the SVD, but the most common approaches actually require computing the matrix-sign function as a subroutine [33, 35]. Numerical analysts have spent decades developing iterative methods for computing $\operatorname*{polar}(\bm{M})$. This rich line of work includes Newton–Schulz [19, Chapter 8], Padé iteration [25, 18], the Newton and scaled Newton iterations [19, Chapter 8], the QWHD iteration [32, 35], and Zolo-pd (Zolotarev polar decomposition) [33]. Unfortunately, as discussed in Section 2, most of these methods are based on rational approximations to the function $\operatorname*{sign}(x)$ and require computing matrix inverses or QR decompositions. Such methods are ill-suited to GPU acceleration and deep learning applications. In contrast, the older Newton-Schulz method is based on polynomial approximation of $\operatorname*{sign}(x)$ and uses only matrix-matrix products. Thus, Muon initially used Newton-Schulz [4]. Indeed, Muon stands for “MomentUm Orthogonalized by Newton-Schulz” [22].
@@ 1.3   Contributions
We present Polar Express, an iterative method for approximating $\operatorname*{polar}(\bm{M})$. Our method dynamically adapts the polynomial update rule at each iteration, prioritizing rapid progress in the initial stage and high accuracy in the later stage. Polar Express constructs polynomials $p_{1},\ldots,p_{T}$ so that the resulting composition is the optimal approximation to the sign function with respect to the supremum ($L^{\infty}$) norm (Theorem 4.1). By iteratively applying these polynomials to $\bm{M}$, Polar Express computes an approximation to $\operatorname*{polar}(\bm{M})$ that is optimal in the worst-case at every iteration. Our method converges to $\operatorname*{polar}(\bm{M})$ super-exponentially (Theorem 4.3), and it quickly reaches a good approximation within just five to ten iterations. This early-stage acceleration is especially valuable in deep learning applications, where runtime efficiency takes precedence over high accuracy. In contrast, classical methods like Newton-Schulz suffer from a slow initial convergence, while recent heuristic proposals [22, 9] fail to converge. Our method is efficient to run on GPUs, using only a few matrix-matrix products per iteration. [^1] In Appendices E and F, we describe two further algorithmic ideas that can be incorporated into Polar Express. They are not used in our Muon experiments but they may be beneficial in other settings, and we believe they merit further study.
@@ 2   Related Work
The polynomials used in Padé methods are chosen to match the value and first few derivatives of $\operatorname*{sign}(x)$ at the points $x=\pm 1$. For instance, the update rule of the third method in this family is defined by $p(x)=\frac{1}{16}\left(35x-35x^{3}+21x^{5}-5x^{7}\right)$, which is the unique degree-7 polynomial satisfying $p(\pm 1)=\pm 1$ and $p^{\prime}(\pm 1)=p^{\prime\prime}(\pm 1)=p^{\prime\prime\prime}(\pm 1)=0$. These methods converge so long as all singular values of $\bm{X}_{0}$ lie in $(0,1]$, a condition guaranteed by the initialization of (2). Furthermore, the order of convergence of the degree $12q+$ 1 method is $q+1$ [6]. In particular, the Newton-Schulz method ($q=1$) converges quadratically.
@@
Newton’s method also converges quadratically. Like Newton-Schulz, it works because the rational function $r(x)=\frac{1}{2}(x+x^{-1})$ has a stable fixed point at $1$; unlike for Newton-Schulz, this point is a global attractor for the whole positive real line. At first glance, Newton’s method has nothing to do with the Padé iterations discussed above. However, after a change of variables $\bm{Y}_{t}=\bm{X}_{t}^{-1}$, it can be reinterpreted as $\bm{Y}_{t+1}=2\bm{Y}_{t}(\bm{I}+\bm{Y}_{t}^{\top}\bm{Y}_{t})^{-1}$, which is sometimes called inverse Newton. Observing that $r(x)=\frac{2x}{1+x^{2}}$ satisfies $r(\pm 1)=\pm 1$ and $r^{\prime}(\pm 1)=0$, we see that (inverse) Newton is also a Padé method, though a rational rather than polynomial one. In fact, given a odd degree $12q_{n}+$ 1 for the numerator and an even degree $2q_{d}$ for the denominator, there is a unique rational function that matches the value and first $q_{n}+q_{d}$ derivatives of $\operatorname*{sign}(x)$ at $x=\pm 1$. This directly yields a Padé method for computing $\operatorname*{polar}(\bm{M})$ whose order of convergence is $q_{n}+q_{d}+1$. For instance, $r(x)=\frac{3x+x^{3}}{1+3x^{2}}$ is called Halley’s method, which converges cubically. When $q_{d}=0$, we recover the polynomial Padé methods.
@@
A landmark 2016 paper introduced a new paradigm to design iterative methods for computing $\operatorname*{polar}(\bm{M})$ [33]. We describe this paradigm in more detail in Section 4, but the main insight is as follows. Padé methods choose the update rule to be an approximation to $\operatorname*{sign}(x)$ of a given degree that is optimally accurate in the neighborhood of $x=1$. Instead, we should choose the approximation to $\operatorname*{sign}(x)$ that is optimal over an interval $[\ell,1]\subset\mathbb{R}_{\geq 0}$ that contains the singular values. Moreover, after each step of the algorithm, the range of the singular values changes; therefore, we adapt the update rule at each iteration to match the new interval. When the range of the singular values is large, this approach ensures that the update rule shrinks it as quickly as possible. As the algorithm proceeds and the interval shrinks to a small neighborhood of $1$, the update rule approaches that of a Padé method, maintaining the same high order of convergence as it has.
Within the class of odd rational functions whose numerators and denominators have degree $12q+$ 1 and $2q$, respectively, an explicit formula for this optimal approximation to $\operatorname*{sign}(x)$ on any interval $[\ell,1]$ was found by Zolotarev. It was shown that these rationals have remarkable convergence properties for any $q$ [33]. For $q=1$, this optimal approximation coincides exactly with the dynamically weighted Halley’s method (QDWH) referenced above. For even faster convergence than QDWH, [33] proposed the zolo-pd method, which uses $q=17$. Finally, these methods all admit the same QR-based implementation trick as QDWH.
@@
In this paper, we adopt the paradigm of zolo-pd [33] but with polynomials rather than rationals of degree $(2q+1,2q)$. This choice avoids the need for QR factorizations, relying solely on GPU-friendly matrix-matrix multiplications in low-precision arithmetic. While this class of methods has not been fully developed in the numerical analysis literature, similar ideas have been rediscovered in different guises. In an unpublished manuscript that predates zolo-pd, Chen and Chow [11] describe a rescaling strategy for Newton-Schulz. Though motivated differently, their method is equivalent to ours for degree-3 polynomials. They also observe numerical instability that prevents the method from converging to all the way to machine precision. Using the insights of [34], they propose a simple mitigation for this issue that we adopt in Section 4.4. Our work gives this method a theoretical foundation that connects it to the paradigm of zolo-pd, and we prove its optimality in the sense of (6). In addition, we study odd polynomials of arbitrary degree and focus particularly on the degree-5 case. Independently, a group of cryptographers developed a similar method for approximating the scalar function $\operatorname*{sign}(x)$ in the context of homomorphic encryption schemes [28]. Their focus is mainly on tuning the analogues in their setting of the polynomial degree and number of iterations, whereas we focus on demonstrating optimality and efficiently constructing the update polynomials for degree $3$ and $5$. In addition, we consider matrix-valued inputs in low-precision arithmetic—not scalars in exact arithmetic—and we demonstrate our method’s effectiveness within the Muon algorithm for training deep neural networks.
@@
The designers of Muon realized that, due to the extreme efficiency requirements and lax accuracy requirements of their setting, rational-based methods from the numerical analysis literature are inapplicable. However, polynomial-based iteration schemes can take full advantage of GPUs because they use only matrix-matrix products in half-precision arithmetic, not inverses or QR decompositions. The preference for speed over accuracy motivates methods that aim to quickly produce coarse approximations, even at the cost of asymptotic convergence. Examples include the proposals of Jordan [22] and You [44, 9], as discussed in Section 1.2. Like Chen and Chow [11], Jordan found that convergence in the initial phase can be accelerated by choosing update rules that have a large derivative near zero, so as to increase the small singular values as much as possible at each iteration. You furthermore chose to use different update rules at each iteration, allowing extra flexibility to tune the trade-off between speed and accuracy. Both used degree-5 polynomials that were found through gradient descent on heuristic objective functions. These proposals were previously compared to Newton-Schultz [^3] Jordan [22] actually compares to $2x-\frac{3}{2}x^{3}+\frac{1}{2}x^{5}$, whereas the true degree-5 Newton-Schulz polynomial is $(15x-10x^{3}+3x^{5})/8$. However, the difference in performance is negligible for the first few iterations., but never to Chen and Chow’s method. We find that our method outperforms them all.

### v3

@@ Abstract
Computing the polar decomposition and the related matrix sign function has been a well-studied problem in numerical analysis for decades. Recently, it has emerged as an important subroutine within the Muon algorithm for training deep neural networks. However, the requirements of this application differ sharply from classical settings: deep learning demands GPU-friendly algorithms that prioritize high throughput over high precision. We introduce Polar Express, a new method for computing the polar decomposition. [^1] https://github.com/NoahAmsel/PolarExpress Like Newton–Schulz and other classical polynomial methods, our approach uses only matrix-matrix multiplications, making it very efficient on GPUs. Inspired by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the update rule at each iteration by solving a minimax optimization problem. We prove that this strategy minimizes error in a worst-case sense, allowing Polar Express to converge as rapidly as possible both in the early iterations and asymptotically. We also address finite-precision issues, making it practical to use in bfloat16. When integrated into the Muon training framework, our method leads to consistent improvements in validation loss when training a GPT-2 model on one billion tokens from the FineWeb dataset, outperforming recent alternatives across a range of learning rates.
@@ 1  Introduction
X = X / (X.norm(dim=(-2, -1), keepdim=True) * 1.01 +1e-7)
@@
Figure 1: Training a GPT-2-Large model (774M params) on 1 billion tokens from the FineWeb dataset [2]. The label muon-<name> refers to implementing Muon using <name> to compute the polar factor. Left: final validation loss across learning rates. Right: validation loss across epochs using the best learning rate. The best learning rate (lrlr) and final validation loss for each method was adamw $(lr=0.0001)$: $4.172$, muon-You $(lr=0.02)$: $3.400$, muon-Jordan $(lr=0.02)$: $3.398$ and muon-PolarExp $(lr=0.02)$: $3.340$.
@@ 1.1  The Muon Method
The Muon optimizer has recently gained popularity for training large language models, often outperforming state-of-the-art adaptive gradient methods like Adam and AdamW [30, 35]. Muon has been used to set records for the NanoGPT speedrun [25], to expand the Pareto frontier of performance versus training FLOPs for large language models [34, 47], and even to train a 32-billion parameter frontier LLM [29].
@@
Whereas standard stochastic gradient descent (SGD) with momentum updates the weight matrix by taking a step in the direction $\bm{M}_{t}$, the Muon method steps in the direction $\operatorname*{polar}(\bm{M}_{t})$, where $\operatorname*{polar}(\bm{M})$ denotes the closest semi-orthogonal matrix to $\bm{M}$ [22, Chapter 8]. Concretely, if $\bm{M}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}$ is the singular value decomposition (SVD) of $\bm{M}$, then
@@
The matrix $\operatorname*{polar}(\bm{M})$ can be seen as a generalization of the matrix sign function to rectangular matrices [3]. Indeed, when $\bm{M}$ is square symmetric with eigendecomposition $\bm{M}=\bm{V}\bm{\Lambda}\bm{V}^{\mathsf{T}}$, $\operatorname*{polar}(\bm{M})$ exactly coincides with the matrix sign function $\operatorname*{sign}(\bm{M})=\bm{V}\operatorname*{sign}(\bm{\Lambda})\bm{V}^{\mathsf{T}}$ [22, Chapter 5]. Equivalently, $\operatorname*{polar}(\bm{M})$ is the left orthogonal factor of the polar decomposition of $\bm{M}$ [22, Chapter 8]. The motivation for Muon is that $\operatorname*{polar}(\bm{M})$ gives the steepest-descent direction with respect to the spectral norm (instead of the Frobenius norm, as in standard SGD).
Recent work [45] shows that Muon can be viewed as a conditional gradient (Frank-Wolfe) method with a trust region defined by the spectral norm. In the same work, the authors also provide a convergence theory for the smooth and non-convex setting, as well as for the stochastic non-convex case. The analysis of Muon was further refined in [46], which proves convergence under a layerwise $(L_{0},L_{1})$-smoothness assumption, in both the stochastic non-convex and stochastic Polyak–Łojasiewicz settings. We also note earlier work that anticipated Muon’s use of the polar factor and its motivation as steepest descent under the spectral norm [8, 9]. We refer the reader to [25] and [5] for further background. In this paper, we take the Muon update rule as given and focus on the problem of efficiently computing the polar decomposition $\operatorname*{polar}(\bm{M})$.
@@ 1.2  Computing the Polar Factor
Although $\operatorname*{polar}(\bm{M})$ can be computed directly via an SVD in $O(\min(mn^{2},nm^{2}))$ time, doing so is prohibitively expensive in deep learning applications, especially as standard SVD algorithms fail to take full advantage of the parallelism available on GPUs. There has been significant work on highly-parallel methods for the SVD, but the most common approaches actually require computing the matrix-sign function as a subroutine [38, 40]. Numerical analysts have spent decades developing iterative methods for computing $\operatorname*{polar}(\bm{M})$. This rich line of work includes Newton–Schulz [22, Chapter 8], Padé iteration [28, 21], the Newton and scaled Newton iterations [22, Chapter 8], the QWHD iteration [37, 40], and Zolo-pd (Zolotarev polar decomposition) [38]. Unfortunately, as discussed in Section 2, most of these methods are based on rational approximations to the function $\operatorname*{sign}(x)$ and require computing matrix inverses or QR decompositions. Such methods are ill-suited to GPU acceleration and deep learning applications. In contrast, the older Newton-Schulz method is based on polynomial approximation of $\operatorname*{sign}(x)$ and uses only matrix-matrix products. Thus, Muon initially used Newton-Schulz [4]. Indeed, Muon stands for “MomentUm Orthogonalized by Newton-Schulz” [25].
@@ 1.3  Contributions
We present Polar Express [^2] https://github.com/NoahAmsel/PolarExpress, an iterative method for approximating $\operatorname*{polar}(\bm{M})$. Our method dynamically adapts the polynomial update rule at each iteration, prioritizing rapid progress in the initial stage and high accuracy in the later stage. Polar Express constructs polynomials $p_{1},\ldots,p_{T}$ so that the resulting composition is the optimal approximation to the sign function with respect to the supremum ($L^{\infty}$) norm (Theorem 4.1). By iteratively applying these polynomials to $\bm{M}$, Polar Express computes an approximation to $\operatorname*{polar}(\bm{M})$ that is optimal in the worst-case at every iteration. Our method converges to $\operatorname*{polar}(\bm{M})$ super-exponentially (Theorem 4.3), and it quickly reaches a good approximation within just five to ten iterations. This early-stage acceleration is especially valuable in deep learning applications, where runtime efficiency takes precedence over high accuracy. In contrast, classical methods like Newton-Schulz suffer from a slow initial convergence, while recent heuristic proposals [25, 10] fail to converge. Our method is efficient to run on GPUs, using only a few matrix-matrix products per iteration. [^3] In Appendices E and F, we describe two further algorithmic ideas that can be incorporated into Polar Express. They are not used in our Muon experiments but they may be beneficial in other settings, and we believe they merit further study.
@@ 2  Related Work
The polynomials used in Padé methods are chosen to match the value and first few derivatives of $\operatorname*{sign}(x)$ at the points $x=\pm 1$. For instance, the update rule of the third method in this family is defined by $p(x)=\frac{1}{16}\left(35x-35x^{3}+21x^{5}-5x^{7}\right)$, which is the unique degree-7 polynomial satisfying $p(\pm 1)=\pm 1$ and $p^{\prime}(\pm 1)=p^{\prime\prime}(\pm 1)=p^{\prime\prime\prime}(\pm 1)=0$. These methods converge so long as all singular values of $\bm{X}_{0}$ lie in $(0,1]$, a condition guaranteed by the initialization of (2). Furthermore, the order of convergence of the degree $2q+1$ method is $q+1$ [6]. In particular, the Newton-Schulz method ($q=1$) converges quadratically.
@@
Newton’s method also converges quadratically. Like Newton-Schulz, it works because the rational function $r(x)=\frac{1}{2}(x+x^{-1})$ has a stable fixed point at $1$; unlike for Newton-Schulz, this point is a global attractor for the whole positive real line. At first glance, Newton’s method has nothing to do with the Padé iterations discussed above. However, after a change of variables $\bm{Y}_{t}=\bm{X}_{t}^{-1}$, it can be reinterpreted as $\bm{Y}_{t+1}=2\bm{Y}_{t}(\bm{I}+\bm{Y}_{t}^{\top}\bm{Y}_{t})^{-1}$, which is sometimes called inverse Newton. Observing that $r(x)=\frac{2x}{1+x^{2}}$ satisfies $r(\pm 1)=\pm 1$ and $r^{\prime}(\pm 1)=0$, we see that (inverse) Newton is also a Padé method, though a rational rather than polynomial one. In fact, given a odd degree $2q_{n}+1$ for the numerator and an even degree $2q_{d}$ for the denominator, there is a unique rational function that matches the value and first $q_{n}+q_{d}$ derivatives of $\operatorname*{sign}(x)$ at $x=\pm 1$. This directly yields a Padé method for computing $\operatorname*{polar}(\bm{M})$ whose order of convergence is $q_{n}+q_{d}+1$. For instance, $r(x)=\frac{3x+x^{3}}{1+3x^{2}}$ is called Halley’s method, which converges cubically. When $q_{d}=0$, we recover the polynomial Padé methods.
@@
A landmark 2016 paper introduced a new paradigm to design iterative methods for computing $\operatorname*{polar}(\bm{M})$ [38]. We describe this paradigm in more detail in Section 4, but the main insight is as follows. Padé methods choose the update rule to be an approximation to $\operatorname*{sign}(x)$ of a given degree that is optimally accurate in the neighborhood of $x=1$. Instead, we should choose the approximation to $\operatorname*{sign}(x)$ that is optimal over an interval $\ell,1]\subset\mathbb{R}_{\geq 0}$ that contains the singular values. Moreover, after each step of the algorithm, the range of the singular values changes; therefore, we adapt the update rule at each iteration to match the new interval. When the range of the singular values is large, this approach ensures that the update rule shrinks it as quickly as possible. As the algorithm proceeds and the interval shrinks to a small neighborhood of $1$, the update rule approaches that of a Padé method, maintaining the same high order of convergence as it has.
Within the class of odd rational functions whose numerators and denominators have degree $2q+1$ and $2q$, respectively, an explicit formula for this optimal approximation to $\operatorname*{sign}(x)$ on any interval $\ell,1]$ was found by Zolotarev. It was shown that these rationals have remarkable convergence properties for any $q$ [38]. For $q=1$, this optimal approximation coincides exactly with the dynamically weighted Halley’s method (QDWH) referenced above. For even faster convergence than QDWH, [38] proposed the Zolo-pd method, which uses $q=17$. Finally, these methods all admit the same QR-based implementation trick as QDWH.
@@
In this paper, we adopt the paradigm of Zolo-pd [38] but with polynomials rather than rationals of degree $(2q+1,2q)$. This choice avoids the need for QR factorizations, relying solely on GPU-friendly matrix-matrix multiplications in low-precision arithmetic. While this class of methods has not been fully developed in the numerical analysis literature, similar ideas have been rediscovered in different guises. In an unpublished manuscript that predates Zolo-pd, Chen and Chow [12] describe a rescaling strategy for Newton-Schulz. Though motivated differently, their method is equivalent to ours for degree-3 polynomials (unlike our work, they do not consider general odd degree). They also observe numerical instability that prevents the method from converging to all the way to machine precision. Using the insights of [39], they propose a simple mitigation for this issue that we adopt in Section 4.4. Our work gives the approach from [39] a stronger theoretical foundation that connects to the paradigm of Zolo-pd. Concretely, we prove that choosing an optimal polynomial at each iteration leads to a composed polynomial that is globally optimal in the sense of (6).
Independently, a group of cryptographers developed a similar method for approximating the scalar function $\operatorname*{sign}(x)$ in the context of homomorphic encryption schemes [32]. Their focus is mainly on tuning the analogues in their setting of the polynomial degree and number of iterations, whereas we focus on demonstrating optimality and efficiently constructing the update polynomials for degree $3$ and $5$. In addition, we consider matrix-valued inputs in low-precision arithmetic—not scalars in exact arithmetic—and we demonstrate our method’s effectiveness within the Muon algorithm for training deep neural networks.
@@
The designers of Muon realized that, due to the extreme efficiency requirements and lax accuracy requirements of their setting, rational-based methods from the numerical analysis literature are inapplicable. However, polynomial-based iteration schemes can take full advantage of GPUs because they use only matrix-matrix products in half-precision arithmetic, not inverses or QR decompositions. The preference for speed over accuracy motivates methods that aim to quickly produce coarse approximations, even at the cost of asymptotic convergence. Examples include the proposals of Jordan [25] and You [36, 10], as discussed in Section 1.2. Like Chen and Chow [12], Jordan found that convergence in the initial phase can be accelerated by choosing update rules that have a large derivative near zero, so as to increase the small singular values as much as possible at each iteration. You furthermore chose to use different update rules at each iteration, allowing extra flexibility to tune the trade-off between speed and accuracy. Both used degree-5 polynomials that were found through gradient descent on heuristic objective functions. These proposals were previously compared to Newton-Schultz [^5] Jordan [25] actually compares to $2x-\frac{3}{2}x^{3}+\frac{1}{2}x^{5}$, whereas the true degree-5 Newton-Schulz polynomial is $(15x-10x^{3}+3x^{5})/8$. However, the difference in performance is negligible for the first few iterations., but never to Chen and Chow’s method from [39]. We find that our method (which generalizes [39]) outperforms them all.
Finally, we remark that concurrent work of Grishina, Smirnov, and Rakhuba also proposes an adaptive polynomial method that generalizes [39] and applies it to accelerating Muon [19]. Like [39], this work does not establish global optimality of the composed polynomial as we do in Section 4 or address finite precision considerations.